# VisEscape: A Benchmark for Evaluating Exploration-driven Decision-making in Virtual Escape Rooms

### v1

@@ Abstract
Escape rooms present a unique cognitive challenge that demands exploration-driven planning: players should actively search their environment, continuously update their knowledge based on new discoveries, and connect disparate clues to determine which elements are relevant to their objectives. Motivated by this, we introduce VisEscape, a benchmark of 20 virtual escape rooms specifically designed to evaluate AI models under these challenging conditions, where success depends not only on solving isolated puzzles but also on iteratively constructing and refining spatial-temporal knowledge of a dynamically changing environment. On VisEscape, we observed that even state-of-the-art multimodal models generally fail to escape the rooms, showing considerable variation in their levels of progress and trajectories. To address this issue, we propose VisEscaper, which effectively integrates Memory, Feedback, and ReAct modules, demonstrating significant improvements by performing 3.7 times more effectively and 5.0 times more efficiently on average.
@@ 1 Introduction
Figure 2: An illustration of a trajectory excerpt from one room in VisEscape. To escape the room successfully, agents must explore multiple directions and diverse views, and interact with various objects. Additionally, they need to infer associations between two or more scenes located in different places to solve creative puzzles.
@@
To evaluate agents under these conditions, we propose VisEscape, a benchmark of 20 virtual escape rooms for assessing exploration-centric problem-solving capacities of multimodal agents. Given the implicit and unspecified goal of “escape the room”, the agents explore their surroundings, identify and integrate relevant information, discover how objects can be used, and solve complex puzzles to successfully escape. Additionally, they need to retain previously observed information and interpret indirect clues, requiring them to reason beyond immediate observations and dynamically adapt to environmental changes.
Additionally, VisEscape provides observation as an image for each state and present executable actions in text form. This brings together the advantages of text-based games, which can define diverse action types (e.g., pouring coffee from a pot or cutting a chain with a nipper), and the benefit of being able to focus on evaluating the agent’s planning or reasoning of diverse VLMs by isolating the execution of actions from low-level action controlling.
On VisEscape, we evaluated various VLMs’ performance and observed that even advanced models like Claude-3.5-Sonnet and GPT-4o showed low escape rates of only 10%, while all other models achieved 0% without memory management, reflective thinking, or reasoning, prompting us to propose VisEscaper framework, a strong baseline with three modules—Memory, Feedback, and ReAct ([45])—enabling effective and efficient problem solving on VisEscape. Equipped with VisEscaper framework, models previously wholly unsuccessful in escape successfully escaped, while models with pre-existing escape capabilities demonstrated $\times$3.7 enhanced effectiveness and $\times$5.0 improved efficiency. Additionally, we discovered that frequent failures of low-performing agents involved repetition of meaningless actions, and that their deficiencies could be mitigated with enhanced reasoning capacity.
In summary, the key points of this paper are three-fold:
@@
  • We addressed the deficiencies on reasoning capacity of VLMs by leveraging the reasoning capacity of LLMs, thereby offering valuable insights into potential improvements for multimodal agents.
@@ 2.1 Interactive environments for evaluating LLM/VLM
Traditionally, text-based games have served as a primary benchmark for language models to explore given environments and learn behaviors in an interactive world [7, 16, 33]. More recently, text-based benchmarks have been used as a venue to evaluate various capabilities and tendencies of LLMs in dynamic situations, such as utilizing scientific knowledge [40], ethical decision-making [31].
@@
Meanwhile, the GUI environment has recently been proposed as a benchmark [9, 8, 41] for evaluating VLM’s interaction with visual content. This benchmark was designed to evaluate VLMs’ ability to perceive visual elements on web interfaces[23, 6], reasoning about their functionality [34], and executing actions to successfully complete given tasks [35, 12, 44, 50].
@@ 2.2 LLM/VLM as an action decision maker
To support these adaptation skills, we integrated three specialized modules into out agent architecture; Memory, Feedback and ReAct [49, 52, 29]. Together, these modules strengthen agents’ reasoning, planning, and reflection capabilities, enabling effective performance in dynamic settings.
@@ 3.1 Key Characteristics of VisEscape
3.1.1 Self-directed successive task
In all rooms of VisEscape, the only instruction is “escape the room”, without any guidance on how to progress. Under this uncertainty about where to begin, the agent should achieve an average of 8.4 checkpoints (critical tasks that act as bottlenecks in achieving goal completion), which requires 27.2 steps on average to escape successfully even for the oracle trajectory. Agents must independently explore, discovering objectives through trial and error. This setup requires them to build and adjust their mental model of the environment, recognizing key objects, spatial relationships, and potential interactions. Success depends on the agent’s ability to hypothesize, test, and adapt their understanding. The task mirrors real-world reasoning challenges, where goals must be inferred, and solutions emerge through engagement, with earlier discoveries becoming relevant later in the process.
3.1.2 Diverse interaction with objects
Since all actions available in VisEscape are defined as text, it was possible to define numerous diverse and creative actions that fit with objects placed in the room. For example, an agent can discover a key hidden inside bread by cutting the bread with a knife, open a closet by cutting chains with nippers, and check a hidden code by flipping an hourglass. These interactions are determined by the type of object currently being viewed and the items currently possessed. When an agent performs an action that satisfy the room’s predefined conditions, it triggers transitions in the room state.
Figure 3: Examples of visual quiz that requires associative thinking in VisEscape. For all problems, we controlled the spatial positioning of the locks and clue images to ensure they remain distinctly separated. This design forces the agent not only to store each piece of information independently but also to identify relevant connections between them (such as “color” in the example above) and reason based on the established associations between these separate information elements.
@@ 3.1.2 Diverse interaction with objects
3.1.3 Associative visual quiz
One key element of escape rooms is finding connections between different information, which requires associative thinking—the ability to link seemingly unrelated elements and a fundamental component of creativity [27, 4, 3]. For example, associative thinking helps players connect a symbol on a puzzle with a similar symbol elsewhere and recognize how objects sharing the same color might be related. Since the escape room is optimal for evaluating the associative thinking ability of VLMs that remained unexplored, we placed at least 1-2 number/letter code locks and corresponding hand-crafted visual quizzes in each room, which require such reasoning. See Figure 3 for examples.
@@ 3.2 Dataset composition
VisEscape consists of 20 distinct rooms. Each room comprises four directional walls, each containing multiple objects that are classified into two types: Receptacles and Items, following the object categorization criteria used in [38]. Receptacles are objects directly located within a room; they can be secured with locks or locking mechanisms and are capable of concealing or containing items. Items can be collected into the inventory, and they serve dual functions: facilitating interaction with other receptacles or providing the necessary clues to unlock the locks. VisEscape has 32 different receptacles and 23 different items, with each room containing a distinctive visual quiz. Each room consists of an average of 5.9 receptacles, 5.1 items, and 1-2 visual quizzes. Statistics and examples are shown in Table 1 and Figure 4.
@@ 3.3 Game graph
We construct a game graph to model the state of each room in VisEscape and provide visual observations corresponding to the current state of the game. The agent can “inspect [object]” to examine receptacles or items more closely by switching to a close-up view, navigate to different walls using “turn to [direction]”, and interact with objects by performing actions defined by each object. Each room has an average of 54.2 unique scenes, which transition depending on the actions performed.
@@ 3.4 Room creation process
To design the room, we utilize three different 3D modeling software-Trimble Sketchup, Autodesk Revit, and Chaos Enscape. Specifically, Trimble Sketchup is utilized for object design, Autodesk Revit for object placement and the overall scene composition, and Chaos Enscape for photorealistic rendering. Refer to the supplementary material for the detailed dataset construction process.
Figure 5: An overview of the three modules that constitute VisEscaper and examples of inputs and outputs for each module.
@@ 4 VisEscaper: Exploring Agent in VisEscape
Figure 5 provides an overview of the VisEscaper framework - comprises 3 modules - Memory module, Feedback module, and ReAct [45] module. Each module, along with task-specific instructions, receives information from the environment and other modules’ responses selectively, allowing it to efficiently perform tasks by only considering the information essential for that specific module. For the Memory module, since directly using images from multiple turns of observations as input can exceed the model’s context-length capacity, we utilized the same VLM as an image captioner to replace visual observations with textual descriptions.
@@ 4.1 Memory module
Since the agent can only observe a limited portion of the room, which causes the entire state of the room to be partially observable, the agent must be equipped with memory capabilities in order to make progress. VisEscaper’s memory consists of 3 component-spatial memory, exploration memory, and action memory.
  • Spatial memory: Following the hierarchical structure of the room state, hierarchically store information about room walls, the receptacles on those walls, the states of these receptacles, along with their visual information and characteristics.
  • Exploration memory: Store information about which objects and items have been inspected so far and which have not been inspected. This encourages the agent to explore unobserved information rather than get stuck with already observed information.
  • Salient action memory: Selectively store actions that are worth remembering, such as interactions with objects, attempting to enter passwords, etc.
Memories of VisEscaper serve a crucial function in mitigating the computational constraints imposed by extended logs during trajectory, and addressing the inherent context window constraints of the model while organizing verbose information into structured and prioritized data.
@@ 4.2 Feedback module
The Feedback module takes the action performed by the agent along with the preceding and subsequent observations, allowing it to independently analyze the outcomes or impacts of its actions. As illustrated in Figure 5, even if the observations remain unchanged, the Feedback module can analyze the action ”Enter $0$” and conclude that ”$0$ is not correct password for this lock.” Such action feedback enables the action module, which makes action decisions through independent inference at each step, to select its current actions based on its past behaviors and their outcomes.
@@ 4.3 ReAct module
The ReAct module makes final action decisions by integrating the current observation with abstract and structured memory information transmitted from other modules. Inspired by previous research [45], the agent engages in a thinking process prior to action execution, facilitating reasoned decision-making based on the comprehensive information provided. In addition the agent transmits its reasoning generated in the previous step to the next step, enabling it to track consistent short-term plans across multiple steps.
@@ 5 Experiments
Claude-3.5-Sonnet VisEscaper 21.7 (x3.42) 32.37 (x2.91) 10.92 (x3.26) 249.5 (x0.87)
BaseAgent 6.33 11.11 3.35 286.7
GPT-4o VisEscaper 13.3 (x4.00) 29.02 (x1.81) 3.91 (x6.51) 276.2 (x0.94)
BaseAgent 3.33 16.00 0.60 294.5
@@
Table 2: Comparison with random baseline, BaseAgent, and VisEscaper for Claude-3.5, GPT-4o, and GPT-4o-mini. Refer to the supplementary material for results of other models. The highest improvement ratios compared to the BaseAgent in VisEscaper on each metric are bolded, and the highest scores of the BaseAgent on each metric are underlined.
@@ 5.1 Experiment setup
For each room, we conducted three different trials and averaged the resulting metrics. To prevent a few challenging puzzles from blocking access to subsequent checkpoints, we experimented with both the default no-hint setting ($exp_{\text{base}}$) and the hint-provided setting ($exp_{\text{hint}}$). In $exp_{\text{hint}}$, if there are no changes in the achieved checkpoints for over 30 consecutive turns, hint guideline for subsequent checkpoint and an additional prompt is given, which guides the agent to follow hint guideline. Furthermore, each experiment was terminated early if no significant changes occurred within 100 consecutive turns, and was automatically concluded upon reaching 300 total steps.
@@
Table 3: The results of diverse VLMs applied to VisEscaper across both $exp_{\text{base}}$ and $exp_{\text{hint}}$. For each experimental setting and metric, The best scores are bolded and the second-best scores are underlined. Experiments without steps all ended in failure due to reaching the maximum turn limit. For human scores, we assigned four rooms per participant, with a total of 20 participants in the $exp_{\text{base}}$setting. Refer to the supplementary material for detailed information on the human study.
@@ 5.2 Metrics
  • Success weighted by Path Length (SPL): Success rate weighted by below term:
@@
For $exp_{\text{hint}}$, We define a metric regarding hint usage of agent: Hint-assisted Completion Rate (HCR): The ratio of checkpoints completed by the agent with the assistance of hints. Higher is better for all metrics except for HCR and Step, where lower is better.
@@ 5.3 Baseline
For comparison, we conducted controlled experiments with a BaseAgent, which ablates all three modules of VisEscaper. This experiment evaluated how much progress could be achieved solely with the recent history of actions and observations, excluding the constituent modules. For BaseAgent, only the most recent 20 observations and actions are alternately provided. We omitted open-source models since the success rate of BaseAgent with these models was 0 across all experiments.
According to Table 2, Claude-3.5-Sonnet, GPT-4o, and GPT-4o-mini all demonstrated significant improvements across all metrics-specifically increase $\times$3.7 in success rate, $\times$2.4 in goal completion, $\times$5.0 in SPL, and 7.3% decrease in total step on average.
Notably, although higher success rates typically require more turns, all models with VisEscaper nevertheless demonstrated a reduction in average step count, leading to SPL exhibiting the most substantial improvement among all metrics. This result underscores that the implementation of VisEscaper has made the exploration and decision-making processes of the original VLM agent more effective and efficient.

### v2

@@ Abstract
Escape rooms present a unique cognitive challenge that demands exploration-driven planning: players should actively search their environment, continuously update their knowledge based on new discoveries, and connect disparate clues to determine which elements are relevant to their objectives. Motivated by this, we introduce VisEscape, a benchmark of 20 virtual escape rooms specifically designed to evaluate AI models under these challenging conditions, where success depends not only on solving isolated puzzles but also on iteratively constructing and refining spatial-temporal knowledge of a dynamically changing environment. On VisEscape, we observe that even state-of-the-art multimodal models generally fail to escape the rooms, showing considerable variation in their levels of progress and trajectories. To address this issue, we propose VisEscaper, which effectively integrates Memory, Feedback, and ReAct modules, demonstrating significant improvements by performing 3.7 times more effectively and 4.9 times more efficiently on average compared to baseline agents.111[Uncaptioned image] Code: pull-ups/VisEscape
@@ 1 Introduction
Figure 2: An illustration of an excerpt of a trajectory from one room in VisEscape. To escape the room successfully, agents must explore multiple directions and diverse views, and interact with various objects. Additionally, they need to infer associations between two or more scenes in different locations to solve creative puzzles.
@@
To evaluate agents under these conditions, we propose VisEscape, a benchmark of 20 virtual escape rooms for assessing exploration-centric problem-solving capacities of multimodal agents. Given the implicit and unspecified goal of “escape the room”, the agents explore their surroundings, identify and integrate relevant information, discover how objects can be used, and solve complex puzzles to successfully escape. Moreover, they need to retain previously observed information and interpret indirect clues, requiring them to reason beyond immediate observations and dynamically adapt to environmental changes.
Additionally, VisEscape provides observation as an image for each state and presents executable actions in text form. This brings together the advantages of text-based games, which can define diverse action types (e.g., pouring coffee from a pot or cutting a chain with a nipper), and the benefit of being able to focus on evaluating the agent’s planning or reasoning capabilities of diverse VLMs by isolating the execution of actions from low-level action control.
On VisEscape, we evaluate various VLMs’ performance and observe that even advanced models like Claude-3.5-Sonnet and GPT-4o show low escape rates of only 10%, while all other models achieve 0% without memory management, reflective thinking, or reasoning, prompting us to propose VisEscaper framework, a strong baseline with three modules—Memory, Feedback, and ReAct [45]—enabling effective and efficient problem solving on VisEscape. Equipped with VisEscaper framework, models previously wholly unsuccessful in escaping successfully escaped, while models with pre-existing escape capabilities demonstrate $\times$3.7 enhanced effectiveness and $\times$4.9 improved efficiency. Additionally, we discover that frequent failures of low-performing agents involve repetition of meaningless actions, and that their deficiencies can be mitigated with enhanced reasoning capacity.
In summary, the key points of this paper are threefold:
@@
  • We address the deficiencies in reasoning capacity of VLMs by leveraging the reasoning capacity of LLMs, thereby offering valuable insights into potential improvements for multimodal agents.
@@ 2.1 Interactive Environments for Evaluating LLM/VLM
Traditionally, text-based games have served as a primary benchmark for language models to explore given environments and learn behaviors in an interactive world [8, 17, 33]. More recently, text-based benchmarks have been used as a venue to evaluate various capabilities and tendencies of LLMs in dynamic situations, such as utilizing scientific knowledge [40] and ethical decision-making [31].
@@
Meanwhile, the GUI environment has recently been proposed as a benchmark [10, 9, 41] for evaluating VLMs’ interaction with visual content. This benchmark was designed to evaluate VLMs’ ability to perceive visual elements on web interfaces[24, 7], reason about their functionality [34], and execute actions to successfully complete given tasks [35, 13, 44, 49].
@@ 2.2 LLM/VLM as an Action Decision Maker
To support these adaptation skills, we integrate three specialized modules into our agent architecture; Memory, Feedback, and ReAct [45]. Together, these modules strengthen agents’ reasoning, planning, and reflection capabilities, enabling effective performance in dynamic settings.
@@ 3.1 Key Characteristics of VisEscape
1. Self-directed successive task. In all rooms of VisEscape, the only instruction is “escape the room”, without any guidance on how to progress. Under this uncertainty about where to begin, the agent should achieve an average of 8.4 checkpoints (critical tasks that act as bottlenecks in achieving goal completion), which requires an average of 27.2 steps to successfully escape, even for the oracle trajectory. Agents must independently explore, discovering objectives through trial and error. This setup requires them to build and adjust their mental model of the environment, recognizing key objects, spatial relationships, and potential interactions. Success depends on the agent’s ability to hypothesize, test, and adapt their understanding. The task mirrors real-world reasoning challenges, where goals must be inferred, and solutions emerge through engagement, with earlier discoveries becoming relevant later in the process.
2. Diverse interaction with objects. Since all actions available in VisEscape are defined as text, it was possible to define numerous diverse and creative actions that fit with objects placed in the room. For example, an agent can discover a key hidden inside bread by cutting the bread with a knife, open a closet by cutting chains with nippers, and check a hidden code by flipping an hourglass. These interactions are determined by the type of object currently being viewed and the items currently possessed. When an agent performs an action that satisfies the room’s predefined conditions, it triggers state transitions within the room.
Figure 3: Examples of visual quizzes that requires associative thinking in VisEscape. For all problems, we controlled the spatial positioning of the locks and clue images to ensure they remain distinctly separated. This design forces the agent not only to store each piece of information independently but also to identify relevant connections between them (such as “color” in the example above) and reason based on the established associations between these separate pieces of information.
@@
3. Associative visual quizzes. One key element of escape rooms is finding connections between different pieces of information, which requires associative thinking—the ability to link seemingly unrelated elements and a fundamental component of creativity [28, 5, 4]. For example, associative thinking helps players connect a symbol on a puzzle with a similar symbol elsewhere and recognize how objects sharing the same color might be related. As escape rooms are ideal for evaluating the associative thinking abilities of VLMs—an area that has remained unexplored—we placed one or two number/letter code locks and corresponding hand-crafted visual quizzes in each room, which require such reasoning. See Figure 3 for examples.
@@ 3.2 Dataset Composition
VisEscape consists of 20 distinct rooms. Each room comprises four walls corresponding to four directions, each containing multiple objects that are classified into two types: Receptacles and Items, following the object categorization criteria used in [38]. Receptacles are objects directly located within a room; they can be secured with locks or locking mechanisms and are capable of concealing or containing items. Items can be collected into the inventory, and they serve dual functions: facilitating interaction with other receptacles or providing the necessary clues to unlock the locks. VisEscape has 32 different receptacles and 23 different items, with each room containing a distinctive visual quiz. Each room contains an average of 5.9 receptacles and 5.1 items, along with one or two visual quizzes. Statistics and examples are shown in Table 1 and Figure 4.
@@ 3.3 Game Graph
We construct a game graph to model the state of each room in VisEscape providing visual observations corresponding to the current state of the game. The agent can “inspect [object]” to examine receptacles or items more closely by switching to a close-up view, navigate around the room by using “turn to [direction]”, and interact with objects by performing actions defined by each object. Each room has an average of 54.2 unique scenes, which change depending on the actions performed.
@@ 3.4 Room Creation Process
To design the room, we utilize three different 3D modeling software programs: Trimble Sketchup, Autodesk Revit, and Chaos Enscape. Specifically, Trimble Sketchup is utilized for object design, Autodesk Revit for object placement and the overall scene composition, and Chaos Enscape for photorealistic rendering. Refer to the Appendix C. for the detailed dataset construction process.
Figure 5: An overview of the three modules comprising VisEscaper, along with examples of inputs and outputs for each module.
@@ 4 VisEscaper: Exploring Agent in VisEscape
Figure 5 provides an overview of the VisEscaper framework, which comprises three modules: the Memory, Feedback, and ReAct [45] modules. Each module, along with task-specific instructions, selectively receives information from the environment and responses from other modules, allowing it to efficiently perform tasks by only considering the information essential for that specific module. For the Memory module, since directly using images from multiple turns of observations as input can exceed the model’s context-length capacity, we utilize the same VLM as an image captioner to replace visual observations with textual descriptions.
@@ 4.1 Memory module
Since the agent can only observe a limited portion of the room, which causes the entire state of the room to be partially observable, the agent must be equipped with memory capabilities in order to make progress. VisEscaper’s memory consists of three components:
  • Structured spatial memory: This memory hierarchically stores information about room walls, the receptacles on those walls, the states of these receptacles, along with their visual information and characteristics. This serves as a directional aid for navigating toward specific receptacles or items.
  • Exploration memory: This memory stores information about which objects and items have been inspected so far and which have not. This encourages the agent to explore unobserved information rather than getting stuck on the information acquired thus far.
  • Salient action memory: This memory selectively stores actions that are worth remembering, such as interactions with objects, attempting to enter passwords, etc. This allows the agent to make action decisions by referencing actions it has performed in the past.
The Memory module of VisEscaper serves a crucial function in mitigating the computational constraints imposed by extended logs during trajectories and addressing the inherent context window constraints of the model while organizing verbose information into structured and prioritized data.
@@ 4.2 Feedback module
The Feedback module takes the action performed by the agent along with the preceding and subsequent observations, allowing it to independently analyze the outcomes or impacts of its actions. As illustrated in Figure 5, even if the observations remain unchanged before and after the action, the Feedback module can conclude that “$0$ is not the correct password for this lock.”, which would be difficult to obtain from a single observation alone. Such feedback enables the agent to determine its current actions based on the analysis of the results of its past actions.
@@ 4.3 ReAct module
The ReAct module makes final action decisions by integrating the current observation with abstract and structured memory information transmitted from other modules. Inspired by previous research [45], the agent engages in a thinking process prior to action execution, facilitating reasoned decision-making based on the comprehensive information provided. In addition, the agent transmits its reasoning from the previous step to the next step, enabling it to maintain consistent short-term plans across multiple steps.
@@ 5 Experiments
Claude-3.5-Sonnet VisEscaper 21.7 (x3.44) 32.37 (x2.91) 10.92 (x3.26) 249.5 (x0.87)
BaseAgent 6.3 11.11 3.35 286.7
GPT-4o VisEscaper 13.3 (x4.03) 29.02 (x1.81) 3.91 (x6.51) 276.2 (x0.94)
BaseAgent 3.3 16.00 0.60 294.5
@@
Table 2: Comparison of random baseline, BaseAgent, and VisEscaper for Claude-3.5, GPT-4o, and GPT-4o-mini. The highest improvement ratios achieved by VisEscaper compared to BaseAgent for each metric are bolded, and the highest scores of the BaseAgent on each metric are underlined.
@@ 5.1 Experiment Setup
For each room, we conducted three different trials and averaged the resulting metrics. To prevent a few challenging puzzles from blocking access to subsequent checkpoints, we experimented with both the default no-hint setting ($exp_{\text{base}}$) and the hint-provided setting ($exp_{\text{hint}}$). In $exp_{\text{hint}}$, if there are no changes in achieved checkpoints for over 30 consecutive turns, a hint guideline message is given to guide the agent toward the next checkpoint. Furthermore, each experiment was terminated early if no significant changes occurred within 100 consecutive turns, and was automatically concluded after a total of 300 steps.
@@
Table 3: The results of diverse VLMs applied to VisEscaper across both $exp_{\text{base}}$ and $exp_{\text{hint}}$. For each experimental setting and metric, the best scores are bolded and the second-best scores are underlined. Experiments without step counts all ended in failure due to reaching the maximum turn limit. For human scores, we assigned four rooms per participant, with a total of 20 participants in the $exp_{\text{base}}$ setting. Refer to Appendix D. for detailed information on the human study.
@@ 5.2 Metrics
  • Success weighted by Path Length [2] (SPL): Success rate weighted by below term:
@@
    This term penalizes the success rate based on the total number of steps taken.
For $exp_{\text{hint}}$, we define a metric regarding the agent’s hint usage: Hint-assisted Completion Rate (HCR), which denotes the ratio of checkpoints completed by the agent with the assistance of hints. Higher is better for all metrics except for HCR and Step, where lower is better.
@@ 5.3 Baseline
For comparison, we conducted controlled experiments with a BaseAgent, which ablates all three modules of VisEscaper on $exp_{\text{base}}$ setting. This experiment evaluated how much progress could be achieved solely with the recent history of actions and observations, excluding the constituent modules. For BaseAgent, only the most recent 20 observations and actions are alternately provided. We omitted open-source models since the success rate of BaseAgent with these models was 0 across all experiments.
According to Table 2, Claude-3.5-Sonnet, GPT-4o, and GPT-4o-mini all demonstrated significant improvements across all metrics-specifically increase $\times$3.7 in success rate, $\times$2.4 in goal completion, $\times$4.9 in SPL, and 7.3% decrease in total steps on average.
Notably, although higher success rates typically require more turns, all models with VisEscaper nevertheless demonstrated a reduction in the average step count, leading to SPL exhibiting the most substantial improvement among all metrics. This result underscores that the implementation of VisEscaper has made the exploration and decision-making processes of the original VLM agent both more effective and efficient.